OpenStudy (anonymous):

If 5x - 3y = 7 and 2y - 4x = 3, then 2x - 2y =

OpenStudy (anonymous):

i have the answer it's 20 but I need to know the steps

OpenStudy (texaschic101):

5x - 3y = 7 ===> (4) -4x + 2y = 3 ===>(5) ----------------- 20x - 12y = 28 (result of multiplying by 4) -20x + 10y = 15 (result of multiplying by 5) ----------------add -2y = 43 y = - 43/2 -4x + 2y = 3 -4x + 2(-43/2) = 3 -4x - 86/2 = 3 -4x - 43 = 3 -4x = 3 + 43 -4x = 46 x = - 46/4 which reduces to - 23/2 5x - 3y = 7 5(-23/2) - 3y = 7 - 115/2 - 3y = 7 -3y = 7 + 115/2 -3y = 14/2 + 115/2 -3y = 129/2 y = (129/2) / -3 y = 129/2 * -1/3 y = -129/6 which reduces to - 43/2 lets check our answers before we go any further 5x - 3y = 7 5(-23/2) - 3(-43/2) = 7 - 115/2 + 129/2 = 7 14/2 = 7 7 = 7 (correct) so we know that x = -23/2 and y = -43/2 2x - 2y = 2(-23/2) - 2(-43/2) = - 46/2 + 86/2 = 40/2 = 20

OpenStudy (anonymous):

hey why didn't u just use the system of equations?

OpenStudy (texaschic101):

I used the system of equations and solved for x and y...through elimination

OpenStudy (texaschic101):

you could also use the system of equations and solve for x and y through substitution

OpenStudy (texaschic101):

either way, you are going to get fractions as x and y.

OpenStudy (anonymous):

damn.... okay i got it now thank you so much